i need the answer because i have and they are both no solotution -2x+5x-9=3(x-4)-5
Answer by Alan3354(69443)

i need the answer because i have and they are both no solotution -2x+5x-9=3(x-4)-5
-------------
-2x+5x-9=3(x-4)-5
3x - 9 = 3x - 12 - 5
3x - 9 = 3x - 17
-9 = -17
That's not true --> inconsistent.
--> no solution.
There's no solution because no value of x will make -9 = -17.
If you graph the 2 equations, the lines will be parallel.
They never cross, no intersection, no solution.